OpenStudy (anonymous):

Find all numbers that are not domain of the function. Then give the domain in set notation.

OpenStudy (anonymous):

\[f(x)=\left(\begin{matrix}2x-2 \\ 2x+1\end{matrix}\right)\]

OpenStudy (anonymous):

x cannot be \[-\frac{1}{2}\] otherwise the denominator would be zero and you cannot divide by zero

OpenStudy (anonymous):

so there are no excluded values?